OpenStudy (anonymous):

find y- intercept of the following equation. 3x + y = 10

OpenStudy (anonymous):

put the equation into point slope form y = mx + b b is the y intercept

OpenStudy (anonymous):

sorry, I mean put into linear equation

OpenStudy (anonymous):

another way to find is to solve for y when x = 0

OpenStudy (the_fizicx99):

Set x for 0 and solve for y or put it in y = mx + b 3x + y = 10 -3x -3x y = -3x + 10

OpenStudy (bhandarialex66):

The y-intercept for the above equation is 10.